# LLM2: Let Large Language Models Harness System 2 Reasoning

### v1

@@ Abstract
LLM2: Let Large Language Models Harness System 2 Reasoning
Cheng Yang1,2 thanks: Equal Contribution. This work was completed during Cheng Yang’s time at Tsinghua University.  Chufan Shi2∗  Siheng Li1∗  Bo Shui2  Yujiu Yang2  Wai Lam1 1The Chinese University of Hong Kong 2Tsinghua University yangc21@mails.tsinghua.edu.cn Correspondence: sihengli24@gmail.comyang.yujiu@sz.tsinghua.edu.cn
@@ 1 Introduction
In this paper, we introduce LLM2, which aims to empower LLMs with System 2 reasoning. As shown in Figure 1, LLM2 integrates an LLM (System 1) with a process-based verifier (System 2). During inference, the LLM generates multiple candidates at each time step, and the verifier provides timely feedback on each candidate. By efficiently exploring the generation space based on the verifier’s feedback, LLM2 ultimately identifies more effective outputs. During the training stage, the process-based verifier is optimized with a pairwise comparison loss to distinguish between desirable and undesirable tokens. To obtain informative token pairs data for process-supervision, we propose a token quality exploration strategy that generate synthetic data based on the potential impact of tokens on the generated text.
We evaluate LLM2 on two representative mathematical reasoning datasets: GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021). With the integration of System 2 reasoning, LLM2 achieves substantial performance improvement across Llama3 models ranging from 1B to 8B parameters. For instance, compared to the vanilla Llama3-1B, LLM2 significantly improves accuracy from 50.3 to 57.8 (+7.5) on GSM8K, and from 24.2 to 28.8 (+4.6) on MATH. Combining LLM2 with self-consistency further boosts the model’s performance, enhancing major@20 accuracy from 56.2 to 70.2 (+14.0) on GSM8K. Further analysis on the utilization of self-generated answers underscores the effectiveness and promising potential of synthetic process-supervision data.
@@ 2.1 Dual-process LLM
We aim to build a dual-process LLM (i.e., LLM2), where an LLM serves as System 1 for giving plausible proposals and a verifier functions as System 2 for deliberate thinking, to refine and prevent mistakes introduced by System 1. Specifically, we formalize this procedure as:
@@ 2.3 Synthetic Process-supervision
To create $x_{t}^{-}$, one can sample tokens from the distributions predicted by LLMs. However, LLMs may assign high probability to alternative correct tokens, which leads to false $x_{t}^{-}$ and confuses the training of the verifier. To alleviate this issue, we introduce a token quality exploration strategy for sampling $x_{t}^{-}$. Specifically, the token quality exploration strategy evaluates the quality of individual tokens based on their potential impact on the generated text. This strategy involves three key steps:
@@
The token quality exploration strategy enables the identification of tokens likely to lead to low-quality outputs, providing informative negative examples for training the verifier. In this work, we consider the top-$k$ most probable tokens according to the LLM’s distribution as candidate set, which reduces the computational cost while still captures the most relevant candidates for $x_{t}^{-}$.
@@ 3.1 Experimental Setup
Our experiments are based on the Llama3 model series, specifically using 1B, 3B and 8B instruct versions (Dubey et al., 2024). We leverage these LLMs as the System 1, and utilize them to initialize corresponding verifiers. We use the GSM8K training set as $\mathcal{D}$, and employ the LLMs to generate corresponding synthetic datasets $\mathcal{D}_{s}$ for training verifiers. For evaluation, we utilize two benchmarks: GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). Further details regarding our experimental setup can be found in Appendix A.
@@ 4.1 Self-generated Answers for Synthetic Process-supervision
We further refine our methodology by utilizing the model’s self-generated correct answers as $\mathcal{D}$, replacing traditional golden solutions to formulate $\mathcal{D}_{s}$ for training verifiers. Instances that remain uncorrect after multiple samplings are excluded. Our experiments with Llama3-1B, as illustrated in Table 1 indicate that crafting $\mathcal{D}$ from self-generated data enhances the efficacy of LLM2. On GSM8K, performance heightens from 57.8 to 59.7, marking an improvement of 9.4 over the vanilla model. On MATH, results improve from 28.8 to 30.2, signifying a 6.0 increase over the baseline.
@@ 4.3 Latency
We assess the impact of LLM2’s decoding latency and compare with vanilla models on Llama3 model series. Specifically, as shown in Table 2, we report the averaged per-instance inference latency on GSM8K. Since the process-based verifier in LLM2 only performs inference when the LLM provides multiple candidate tokens after the adaptive plausibility constraint, LLM2 introduces an additional 1.21x to 1.25x latency. This latency tends to decrease as the modes’s parameters increase.
@@ 5 Related Work
Training verifiers to explicitly distinguish between desirable and undesirable outputs has been a promising method to improve the capabilities of LLMs. Existing verifier modeling can be broadly classified into two categories: (1) Outcome-based modeling (Shen et al., 2021; Cobbe et al., 2021), which train verifiers to learn how to distinguish between correct and wrong outputs and selects more optimal ones from a number of candidates at inference time. (2) Process-based modeling (Uesato et al., 2022; Lightman et al., 2023), which supervises each reasoning step of the generation process. To alleviate the reliance on human-annotated process-supervision data, Wang et al. (2024) propose to automatically construct process-supervision data, where the correctness of a mathematical reasoning step is defined as its potential to reach the final answer correctly.
@@
LLM2 integrates System 2 during the inference-stage. Specifically, LLM2 leverages a process-based verifier as the System 2 to provide real-time feedback at each token generation step without generating auxiliary content.
@@ 6 Conclusion
While LLM2 demonstrates significant improvements in mathematical reasoning tasks, our exploration does not extend to other reasoning domains such as commonsense reasoning and code generation, due to computational resource constraints. We are optimistic about the potential of LLM2 to generalize well to these additional tasks. However, applying LLM2 to open-ended tasks, like creative writing, presents challenges due to the lack of definitive supervisory signals for synthetic process-supervision. Addressing these challenges offers a promising direction for future research.
@@ Appendix F Case Study
We present two representative cases from GSM8K using Llama3-1B to demonstrates how LLM2 improves mathematical reasoning in Table 10 and 11.

### v2

# Cheng Yang1,2   Chufan Shi2†  Siheng Li1†  Bo Shui2  Yujiu Yang2  Wai Lam1

@@ Abstract
LLM2: Let Large Language Models Harness System 2 Reasoning thanks: The work described in this paper is partially supported by a grant from the Research Grant Council of the Hong Kong Special Administrative Region, China (Project Code: 14200620).
Cheng Yang1,2 thanks: Equal Contribution. This paper was completed during Cheng Yang’s time at Tsinghua University.  Chufan Shi2†  Siheng Li1†  Bo Shui2  Yujiu Yang2  Wai Lam1 1The Chinese University of Hong Kong 2Tsinghua University yangc21@mails.tsinghua.edu.cn Correspondence: sihengli24@gmail.comyang.yujiu@sz.tsinghua.edu.cn
@@ 1 Introduction
In this paper, we introduce LLM2, which aims to empower LLMs with System 2 reasoning. As shown in Figure 1, LLM2 integrates an LLM (System 1) with a process-based verifier (System 2). During inference, the LLM generates multiple candidates at each time step, and the verifier provides timely feedback on each candidate. By efficiently exploring the generation space based on the verifier’s feedback, LLM2 ultimately identifies more effective outputs. During the training stage, the process-based verifier is optimized with a pairwise comparison loss to distinguish between desirable and undesirable tokens. To obtain informative token pairs data for process-supervision, we propose a token quality exploration strategy that generates synthetic data based on the potential impact of tokens on the generated text.
We evaluate LLM2 on two representative mathematical reasoning datasets: GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021). With the integration of System 2 reasoning, LLM2 achieves substantial performance improvement across Llama3 models ranging from 1B to 8B parameters. For instance, compared to the vanilla Llama3-1B, LLM2 significantly improves accuracy from 50.3 to 57.8 (+7.5) on GSM8K, and from 24.2 to 28.8 (+4.6) on MATH. Combining LLM2 with self-consistency further boosts the model’s performance, enhancing major@20 accuracy from 56.2 to 70.2 (+14.0) on GSM8K. Further analysis of the utilization of self-generated answers underscores the effectiveness and promising potential of synthetic process-supervision data.
@@ 2.1 Dual-process LLM
We aim to build a dual-process LLM (i.e., LLM2), where an LLM serves as System 1 for giving plausible proposals and a verifier functions as System 2 for deliberate thinking to refine and prevent mistakes introduced by System 1. Specifically, we formalize this procedure as:
@@ 2.3 Synthetic Process-supervision
To create $x_{t}^{-}$, one can sample tokens from the distributions predicted by LLMs. However, LLMs may assign a high probability to alternative correct tokens, which leads to false $x_{t}^{-}$ and confuses the training of the verifier. To alleviate this issue, we introduce a token quality exploration strategy for sampling $x_{t}^{-}$. Specifically, the token quality exploration strategy evaluates the quality of individual tokens based on their potential impact on the generated text. This strategy involves three key steps:
@@
The token quality exploration strategy enables the identification of tokens likely to lead to low-quality outputs, providing informative negative examples for training the verifier. In this work, we consider the top-$k$ most probable tokens according to the LLM’s distribution as a candidate set, which reduces the computational cost while still capturing the most relevant candidates for $x_{t}^{-}$.
@@ 3.1 Experimental Setup
Our experiments are based on the Llama3 model series, specifically using 1B, 3B and 8B instruct versions (Dubey et al., 2024). We leverage these LLMs as System 1 and utilize them to initialize corresponding verifiers. We use the GSM8K training set as $\mathcal{D}$, and employ the LLMs to generate corresponding synthetic datasets $\mathcal{D}_{s}$ for training verifiers. For evaluation, we utilize two benchmarks: GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). Further details regarding our experimental setup can be found in Appendix A.
@@ 4.1 Self-generated Answers for Synthetic Process-supervision
We further refine our methodology by utilizing the model’s self-generated correct answers as $\mathcal{D}$, replacing traditional golden solutions to formulate $\mathcal{D}_{s}$ for training verifiers. Instances that remain incorrect after multiple samplings are excluded. Our experiments with Llama3-1B, as illustrated in Table 1 indicate that crafting $\mathcal{D}$ from self-generated data enhances the efficacy of LLM2. On GSM8K, performance heightens from 57.8 to 59.7, marking an improvement of 9.4 over the vanilla model. On MATH, results improve from 28.8 to 30.2, signifying a 6.0 increase over the baseline.
@@ 4.3 Latency
We assess the impact of LLM2’s decoding latency and compare it with vanilla models on the Llama3 model series. Specifically, as shown in Table 2, we report the averaged per-instance inference latency on GSM8K. Since the process-based verifier in LLM2 only performs inference when the LLM provides multiple candidate tokens after the adaptive plausibility constraint, LLM2 introduces an additional 1.21x to 1.25x latency. This latency tends to decrease as the modes’s parameters increase.
@@ 5 Related Work
Training verifiers to explicitly distinguish between desirable and undesirable outputs has been a promising method to improve the capabilities of LLMs. Existing verifier modeling can be broadly classified into two categories: (1) Outcome-based modeling (Shen et al., 2021; Cobbe et al., 2021), which train verifiers to learn how to distinguish between correct and wrong outputs and selects more optimal ones from a number of candidates at inference time. (2) Process-based modeling (Uesato et al., 2022; Lightman et al., 2023; Zhu et al., 2023), which supervises each reasoning step of the generation process. To alleviate the reliance on human-annotated process-supervision data, Wang et al. (2024) propose to automatically construct process-supervision data, where the correctness of a mathematical reasoning step is defined as its potential to reach the final answer correctly.
@@
LLM2 integrates System 2 during the inference stage. Specifically, LLM2 leverages a process-based verifier as System 2 to provide real-time feedback at each token generation step without generating auxiliary content.
@@ 6 Conclusion
While LLM2 demonstrates significant improvements in mathematical reasoning tasks, our exploration does not extend to other reasoning domains, such as commonsense reasoning and code generation, due to computational resource constraints. We are optimistic about the potential of LLM2 to generalize well to these additional tasks. However, applying LLM2 to open-ended tasks, like creative writing, presents challenges due to the lack of definitive supervisory signals for synthetic process-supervision. Addressing these challenges offers a promising direction for future research.
Acknowledgments
This work was partly supported by the National Key Research and Development Program of China (No. 2024YFB2808903), the research grant No. CT20240905126002 of the Doubao Large Model Fund and the Shenzhen Science and Technology Program JSGG20220831110203007).
@@ Appendix F Case Study
We present two representative cases from GSM8K using Llama3-1B to demonstrate how LLM2 improves mathematical reasoning in Table 10 and 11.